Solve the the following equations by factoring. Show steps!!
1. 49d^3-81d=0
2. 2x^2=9x+5
3. (2x-3)(x+4)=0
SHOW STEPS

Answer by sachi(548)

1. 49d^3-81d=0
or d(49d^2-81)=0
or d(7d-9)(7d+9)=0
or d=0
or 7d-9=0 =>d=9/7
or 7d+9=0 =>d=-9/7
2. 2x^2=9x+5
or 2x^2-9x-5=0
or 2x^2-10x+x-5=0
or 2x(x-5)+1(x-5)=0
or (x-5)(2x+1)=0
or x-5=0 => x=5
or 2x+1=0 =>x=-1/2
3. (2x-3)(x+4)=0
or 2x-3=0 =>x=3/2
or x+4=0 =>x=-4
ans